OpenStudy (anonymous):

Which of the following best describes the relationship between (x + 1) and the polynomial x2 - x - 2? A. (x + 1) is a factor. B. (x + 1) is not a factor. C. It is impossible to tell whether (x + 1) is a factor.

OpenStudy (welshfella):

can you factor x^2 - x - 2?

OpenStudy (welshfella):

it comes out like (x + a)(x + b) though there might be negatives as well

OpenStudy (welshfella):

to find a and b you need 2 numbers whose product is -2 and whose sum is -1

OpenStudy (anonymous):

i have no idea how to do this at all

OpenStudy (welshfella):

what 2 numbers fit a*b = -2 and a + b = -1?

OpenStudy (welshfella):

OK if you cant do then use the Factor Theorem if (X + 1) is a factor then f(-1) = 0 f(x) = x^2 - x - 2 f(-1) = (-1)^2 - (-1) - 2 - does that work out to 0?

OpenStudy (welshfella):

f(-1) just means f(x) where x = -1

OpenStudy (welshfella):

no - try again (-1)^2 = 1 -(-1) = what?

OpenStudy (anonymous):

wait i think i get it

OpenStudy (anonymous):

would the answer happen to be (x + 1) is not a factor.

OpenStudy (welshfella):

NO (-1)^2 -(-1) - 2 = 1 + 1 - 2 = 0 THEREFORE BY THE FACTOR THEOREM (X + 1) is A FACTOR

OpenStudy (anonymous):

Crap sorry im not good at algebra

OpenStudy (welshfella):

double negative -(-1) = +1 also -1 * -1 = + 1

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

learn the basics and practice some and you'd be surprised...

OpenStudy (welshfella):

gtg

OpenStudy (daniel.ohearn1):

How would you factor that polynomial?

OpenStudy (daniel.ohearn1):

you know the third term in the product you will get is -2

OpenStudy (daniel.ohearn1):

y=x^2+(a+b)x+ab where we have (x+a)(x+b)